OpenStudy (ineedserioushelp):

5. Use the given property to complete the statement. Multiplication Property of Equality If 6x ÷ 8 = 40, then ____. (1 point) 6x * 8 = 320 40 = 6x * 8 <--- My answer 6x = 320 40 = 6x ÷ 8

OpenStudy (ineedserioushelp):

check my answer please ?

OpenStudy (anonymous):

Okay, so when you use the multiplication property of equality, you do it to both sides of the equation.

OpenStudy (anonymous):

6x / 8 = 40 6x /8 * 8 = 40 * 8 so what would that equal?

OpenStudy (ineedserioushelp):

6x = 320?

OpenStudy (ineedserioushelp):

@raimj512

OpenStudy (anonymous):

Yeah, exactly =)

OpenStudy (ineedserioushelp):

thanks =D